# Avoiding strict saddle points of nonconvex regularized problems

### v2

@@ 4 Convergence analysis
$\|f(x)-f(y)\|\leq L_{f}\|x-y\|\text{ \ and \ }\|\nabla f(x)-\nabla f(y)\|\leq L_{f}\|x-y\|.$ (17)
We first show the monotonicity of $F(x,\epsilon)$ along a direction. Define the reductions of $G$ and $F$ caused by $x^{k+1}$ as
@@
$r(|x_{i}^{k+1}|+(\epsilon_{i}^{k})^{2})\leq r(|x_{i}^{k}|+(\epsilon_{i}^{k})^{2})+r^{\prime}(|x_{i}^{k}|+(\epsilon^{k}_{i})^{2})(|x_{i}^{k+1}|-|x_{i}^{k}|).$
@@
$\sum\limits_{i=1}^{n}r(|x_{i}^{k+1}|+(\epsilon_{i}^{k})^{2})-\sum\limits_{i=1}^{n}r(|x_{i}^{k}|+(\epsilon_{i}^{k})^{2})\leq\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon^{k}_{i})(|x_{i}^{k+1}|-|x_{i}^{k}|).$ (19)
@@
(Global convergence) Suppose Assumption 1 and 3 hold. Let $\{(x^{k},\epsilon^{k})\}$ being generated by ${\rm DIRL}_{1}$ and $\Omega$ be the set of the limit points of $\{x^{k}\}$. Then $\mu^{*}=0$ and $\Omega\neq\emptyset$ and any $x^{*}\in\Omega$ is first-order optimal for (P).
@@
Then we consider the two kinds of cases with $r^{\prime}(|0^{+}|)<\infty$ and $r^{\prime}(|0^{+}|)=\infty$.
If $r^{\prime}(|0^{+}|)=\infty$, as shown in Proposition 2, we only need to consider $i\in\mathcal{I}^{*}$. Taking the limit on $\mathcal{S}$,
00 0 $=\lim\limits_{k\rightarrow\infty\atop k\in\mathcal{S}}|\nabla_{i}f(x^{k})+\lambda w(x_{i}^{k},\epsilon_{i}^{k}){\rm sign}(y_{i}^{k})+\beta(y_{i}^{k}-x_{i}^{k})|$
$=\lim\limits_{k\rightarrow\infty\atop k\in\mathcal{S}}|\nabla_{i}f(x^{k})+\lambda w(x_{i}^{k},\epsilon_{i}^{k}){\rm sign}(x_{i}^{k})+\frac{\beta}{\alpha}(x_{i}^{k+1}-x_{i}^{k})|$
$\geq\lim\limits_{k\rightarrow\infty\atop k\in\mathcal{S}}|\nabla_{i}f(x^{k})+\lambda w(x_{i}^{k},\epsilon_{i}^{k}){\rm sign}(x_{i}^{k})|-\frac{\beta}{\alpha}\|x^{k}-x^{k+1}\|$
$=|\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0){\rm sign}(x_{i}^{*})|$
$=|\nabla_{i}f(x^{*})+\lambda r^{\prime}(|x_{i}^{*}|){\rm sign}(x_{i}^{*})|.$
which satisfy (5). Therefore, $x^{*}$ is stationary.
If $r^{\prime}(|0^{+}|)<\infty$, we need also to consider $i\in\mathcal{J}^{*}$. By [41, Proposition 2.1.5] and [37, Theorem 5.19], $\partial r$ is outer semi-continuous and there exists $\xi^{*}_{i}\in\partial|x^{*}_{i}|$ such that
$0=\nabla_{i}f(x^{*})+\lambda w(0,0)\xi^{*}_{i}=\nabla_{i}f(x^{*})+\lambda r^{\prime}(|0^{+}|)\text{sign}(x_{i}^{*}),$
meaning (5) is satisfied. Therefore, $x^{*}$ is stationary. ∎
5 Avoidance of strict saddle point
DIRL1 has the property of so-called “active manifold identification”, which belongs to the research line of active-manifold identification in nonsmooth optimization [42, 43]. It means the iterates fall onto a smooth manifold after finite iterations. The iteratively reweighted $\ell_{1}$ algorithms were proved to have this property by [33, 32] and the manifold in this case is simply the subspace consists of the nonzeros at a limit point.
@@ 5 Avoidance of strict saddle point
(Model Identification and Local Smoothness) Under Assumption 1-3, for any critical point $(x^{*},\epsilon^{*})$, there exists a sufficiently small neighborhood $B$ of $(x^{*},\epsilon^{*})$, so that the following hold:
@@
    For any point $(x,\epsilon)\in B$, $S(x,\epsilon)\in(x^{*}+\mathcal{M}_{\mathcal{I}^{*}})$ and $\text{sign}(S(x,\epsilon))=\text{sign}(x)=\text{sign}(x^{*})$.
@@
    Map $S$ is $C^{1}$-smooth in $B$ and $S(x)=x-\frac{1}{\beta}g(\theta)$ with
    $g_{i}(\theta)=\left\{\begin{aligned} &\beta x_{i},&i\in\mathcal{J}^{*},\\ &\nabla_{i}f(x)+\lambda r^{\prime}(x_{i}\text{sign}(x_{i}^{*})+\epsilon_{i}^{2})\text{sign}(x_{i}^{*}),&i\in\mathcal{I}^{*}.\end{aligned}\right.$
@@
(i) Since $S$ is a continuous map and $x^{*}=S(x^{*},\epsilon^{*})$, it holds that
@@
$S(x^{k},\epsilon^{k})=$ $\ \arg\min\limits_{x\in x^{*}+\mathcal{M}_{\mathcal{I}^{*}}}\{\nabla f(x^{k})^{T}x+\tfrac{\beta}{2}\|x-x^{k}\|^{2}+\lambda\sum_{i=1}^{n}w^{k}_{i}\text{sign}(x_{i}^{*})x_{i}\}$ (24)
@@
The above theorem shows the local behavior of ${\rm DIRL}_{1}$: it is equivalent to the gradient descent for the variables in $\mathcal{I}^{*}$ locally around a critical point in $x^{*}+\mathcal{M}_{\mathcal{I}^{*}}$. Specifically, it can be expressed as
$T_{i}(x,\epsilon)=\left\{\begin{aligned} &(1-\alpha)x_{i},&&i\in\mathcal{J}^{*},\\ &x_{i}-\frac{\alpha}{\beta}\nabla_{x_{i}}F(x,\epsilon),&&i\in\mathcal{I}^{*},\\ &\mu\epsilon_{i},&&i\in\{n+1,\ldots,2n\}\end{aligned}\right.$ (25)
@@
(Unstable Fixed Point) Suppose $x^{*}$ is a critical point of (P). $x^{*}$ is a strict saddle point of (P) if and only if the corresponding $(x^{*},\epsilon^{*})$ with $\epsilon^{*}=0$ is an unstable fixed point of $T$.
@@
First of all, we calculate the Jabobian of $T$ in the neighborhood $B$ described in Theorem 8. It holds that
$\frac{\partial g}{\partial\epsilon_{i}}(x,\epsilon)=\begin{cases}0&\text{ \ if}i\in\mathcal{J}^{*},\\ \lambda r^{\prime\prime}(x_{i}\text{sign}(x_{i}^{*})+\epsilon_{i}^{2})\text{sign}(x_{i}^{*})2\epsilon&\text{ \ if }i\in\mathcal{I}^{*}.\end{cases}$
Let $D_{\epsilon}g(x,\epsilon)$ be the Jacobian matrix of $g(\cdot)$ with respect to $\epsilon$. It follows that $D_{\epsilon}g(x^{*},\epsilon^{*})=\bm{0}$ by the fact that $\epsilon_{i}^{*}=0$ and $r^{\prime\prime}(|x_{i}^{*}|)<+\infty$, $i\in\mathcal{I}^{*}$. On the other hand, $D_{x}g(x^{*},\epsilon^{*})=\begin{bmatrix}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})&\nabla^{2}_{\mathcal{I}^{*}\mathcal{J}^{*}}f(x^{*})\\ \bm{0}&\beta I\end{bmatrix}$. Therefore, the Jacobian of $T$ at $(x^{*},\epsilon^{*})$ is given by
$DT^{*}(x^{*},\epsilon^{*})=$ $\ \begin{bmatrix}(1-\alpha)I+\alpha(I-\tfrac{1}{\beta}D_{x}g(x^{*},\epsilon^{*}))\ &D_{\epsilon}g(x^{*},\epsilon^{*})\\ \bm{0}\ &\mu I\end{bmatrix}=\begin{bmatrix}I-\frac{\alpha}{\beta}D_{x}g(x^{*},\epsilon^{*})\ &\bm{0}\\ \bm{0}\ &\mu I\end{bmatrix}$
$=$ $\ \begin{bmatrix}I-\frac{\alpha}{\beta}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})&I-\frac{\alpha}{\beta}\nabla^{2}_{\mathcal{I}^{*}\mathcal{J}^{*}}f(x^{*})&\bm{0}\\ \bm{0}&(1-\alpha)I&\bm{0}\\ \bm{0}&\bm{0}&\mu I\end{bmatrix}$
@@
Suppose Assumption 1-3 hold. Let $L_{r}:=r^{\prime\prime}(|(r^{\prime})^{-1}(C))$ if $r^{\prime}(|0^{+}|)=\infty$, where $C$ defined in Proposition 6, and $L_{r}:=r^{\prime\prime}(|0^{+}|)$ if $r^{\prime}(|0^{+}|)<\infty$. $T$ is a lipeomorphism when $10<\alpha<$ 1, $\beta>\alpha L_{f}/2$ and $\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))<1$.
@@
If $r^{\prime}(|0+|)<\infty$, $w(x_{i},\epsilon_{i})=r^{\prime}(|x_{i}|,\epsilon_{i}^{2})$ is Lipschitz continuous in $\mathcal{L}(F^{0})$.
@@
$u_{i}(x_{i},\epsilon_{i})=\begin{cases}w(x_{i},\epsilon_{i})=r^{\prime}(|x_{i}|+\epsilon_{i}^{2})&\text{ if }r^{\prime}(|0+|)<\infty\\ \min(w(\underline{x},0),w(x_{i},\epsilon_{i}))=\min(r^{\prime}(\underline{x}),r^{\prime}(|x_{i}|+\epsilon_{i}^{2}))&\text{ if }r^{\prime}(|0+|)=\infty.\end{cases}$
@@
$S(\hat{x},\hat{\epsilon})-S(\tilde{x},\tilde{\epsilon})=$ $\ \mathbb{S}_{\lambda\hat{u}/\beta}(\hat{x}-\tfrac{1}{\beta}\nabla f(\hat{x}))-\mathbb{S}_{\lambda\tilde{u}/\beta}(\tilde{x}-\tfrac{1}{\beta}\nabla f(\tilde{x}))$
@@
where the first inequality is by Lemma A.2 and the second is by (17) and (27). Therefore, $\|\alpha(\hat{x}-S(\hat{x},\hat{\epsilon}))-\alpha(\tilde{x}-S(\tilde{x},\tilde{\epsilon}))\|\leq\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))$ holds for any $\hat{y},\tilde{y}\in\mathcal{L}(F^{0})$ and $\hat{\epsilon},\tilde{\epsilon}\in\mathbb{R}^{n}_{+}$.
Notice that $T=\begin{bmatrix}x\\ \epsilon\end{bmatrix}+\alpha(\begin{bmatrix}S(x,\epsilon)\\ \mu\epsilon\end{bmatrix}-\begin{bmatrix}x\\ \epsilon\end{bmatrix})$, Since $\begin{bmatrix}S(x,\epsilon)\\ \mu\epsilon\end{bmatrix}$ is Lipschitz with constant $\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))<1$. Applying [23, Lemma 2.13], we directly derive that $T=I+\alpha(S-I)$ is a lipeomorphism. ∎
@@
(Global Escape) Consider problem (P) with strict saddle property and the ${\rm DIRL}_{1}$ algorithm with Assumption 1-3. Let the condition of Proposition 10 holds. Let $\{(x^{k},\epsilon^{k})\}$ be the sequence generated by the Algorithm 10. Then, the set of initial points attracted by strict saddle points has zero Lebesgue measure. of (P).

### v3

@@ 3.1 Convergence analysis
$|f(x)-f(y)|\leq L_{f}\|x-y\|\text{ \ and \ }\|\nabla f(x)-\nabla f(y)\|\leq L_{f}\|x-y\|.$ (19)
This assumption also implies that if $r^{\prime}(|0^{+}|)=\infty$, then for any stationary point $x^{*}$, $|x_{i}^{*}|\geq(r^{\prime})^{-1}(L_{f}/\lambda),\forall i\in\mathcal{I}^{*}$ by (6). Therefore, we know $r^{\prime\prime}(|x_{i}^{*}|)>0$ is bounded above and this is also the case for $r^{\prime}(|0^{+}|)<\infty$. Overall, we can conclude
$\|\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})\|\leq\rho$ (20)
on all stationary points in $\mathcal{L}(F^{0})$.
We first show the monotonicity of $F(x,\epsilon)$. Define the reductions of $G$ and $F$ caused by $x^{k+1}$ as
@@
$r(|x_{i}^{k+1}|+(\epsilon_{i}^{k})^{2})\leq r(|x_{i}^{k}|+\epsilon_{i}^{k})+r^{\prime}(|x_{i}^{k}|+\epsilon^{k}_{i})(|x_{i}^{k+1}|-|x_{i}^{k}|).$
@@
$\sum\limits_{i=1}^{n}r(|x_{i}^{k+1}|+\epsilon_{i}^{k})-\sum\limits_{i=1}^{n}r(|x_{i}^{k}|+\epsilon_{i}^{k})\leq\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon^{k}_{i})(|x_{i}^{k+1}|-|x_{i}^{k}|).$ (22)
@@
(Global convergence) Suppose Assumption 1 and 3 hold. Let $\{(x^{k},\epsilon^{k})\}$ being generated by ${\rm DIRL}_{1}$ and $\Omega$ be the set of the limit points of $\{x^{k}\}$. Then $\epsilon^{*}=0$ and $\Omega\neq\emptyset$ and any $x^{*}\in\Omega$ is first-order optimal for (P).
@@
Firstly, boundedness of $\mathcal{L}(F^{0})$ from Assumption 3 implies $\Omega\neq\emptyset$. Then let $x^{*}$ be a limit point of $\{x^{k}\}$ with subsequence $\{x^{k}\}_{\mathcal{S}}\rightarrow x^{*}$ and $\{y^{k}\}_{S}\rightarrow x^{*}$ since $y^{k}=\frac{1}{\alpha}(x^{k}-x^{k+1})+x^{k}$. This also implies $\text{sign}(y_{i}^{k})=\text{sign}(x_{i}^{*}),i\in\mathcal{I}^{*}$ for sufficiently large $k\in\mathcal{S}$. The optimality condition of the $k$th subproblem implies
@@
If $w(x_{i}^{*},0)<+\infty$, it follows from [48, Proposition 2.1.5], [42, Theorem 5.19] and the outer semi-continuity of $\partial r$ that there exists $\xi^{*}_{i}\in\partial|x^{*}_{i}|$ such that
$0=\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0)\xi^{*}_{i}$
as $y_{i}^{k}-x_{i}^{k}\overset{\mathcal{S}}{\to}0$, $x_{i}^{k}\overset{\mathcal{S}}{\to}x_{i}^{*}$ and $\epsilon_{i}^{k}\to 0$. Therefore, for $w(x_{i}^{*},0)=r^{\prime}(|x_{i}^{*}|)<\infty,i\in\mathcal{I}^{*}$, (6) is satisfied; for $w(x_{i}^{*},0)=r^{\prime}(|0^{+}|)<\infty,i\in\mathcal{J}^{*}$, (7) is satisfied by Assumption 2. Therefore, $x^{*}$ is stationary. ∎
3.2 Avoidance of strict saddle point
The map $S$ has the property of so-called “active manifold identification”, which means the iterates fall onto a smooth manifold after finite iterations. The iteratively reweighted $\ell_{1}$ algorithms were proved to have this property by [33, 32] and the manifold in this case is simply the subspace consists of the nonzeros at a limit point.
@@ 3.2 Avoidance of strict saddle point
(Model Identification and Local Smoothness) Under Assumption 1-3, for any stationary point $(x^{*},\epsilon^{*})$, there exists a sufficiently small neighborhood $B$ of $(x^{*},\epsilon^{*})$, so that the following hold:
@@
    For any point $(x,\epsilon)\in B$, $S^{x}(x,\epsilon)\in(x^{*}+\mathcal{M}_{\mathcal{I}^{*}})$ and $\text{sign}(S^{x}(x,\epsilon))=\text{sign}(x)=\text{sign}(x^{*})$.
@@
    $S$ is $C^{1}$-smooth in $B$ and $S^{x}(x,\epsilon)=x-\frac{1}{\beta}g(x,\epsilon)$ with
@@
(i) Since $S$ is a continuous map and $x^{*}=S^{x}(x^{*},\epsilon^{*})$, it holds that
@@
$S^{x}(x^{k},\epsilon^{k})=$ $\ \arg\min\limits_{x\in x^{*}+\mathcal{M}_{\mathcal{I}^{*}}}\{\nabla f(x^{k})^{T}x+\tfrac{\beta}{2}\|x-x^{k}\|^{2}+\lambda\sum_{i=1}^{n}w^{k}_{i}\text{sign}(x_{i}^{*})x_{i}\}$ (27)
@@
The above theorem shows the local behavior of ${\rm DIRL}_{1}$: it is equivalent to the gradient descent for the variables in $\mathcal{I}^{*}$ locally around a stationary point in $x^{*}+\mathcal{M}_{\mathcal{I}^{*}}$. Specifically, it can be expressed as
$T_{i}(x,\epsilon)=\left\{\begin{aligned} &(1-\alpha)x_{i},&&i\in\mathcal{J}^{*},\\ &x_{i}-\frac{\alpha}{\beta}\nabla_{x_{i}}F(x,\epsilon),&&i\in\mathcal{I}^{*},\\ &(1-\alpha(1-\mu))\epsilon_{i},&&i\in\{n+1,\ldots,2n\}\end{aligned}\right.$ (28)
@@
(Unstable Fixed Point) Suppose $x^{*}$ is a stationary point of (P). $x^{*}$ is a strict saddle point of (P) if and only if the corresponding $(x^{*},\epsilon^{*})$ with $\epsilon^{*}=0$ is an unstable fixed point of $T$. Moreover, if $\alpha<\frac{\beta}{\rho}$, $DT(x^{*},\epsilon^{*})$ is invertible.
@@
First of all, we calculate the Jabobian of $T$ in the neighborhood $B$ described in Theorem 12. It holds that $D_{x}g(x^{*},\epsilon^{*})=\begin{bmatrix}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})&\nabla^{2}_{\mathcal{I}^{*}\mathcal{J}^{*}}f(x^{*})\\ \bm{0}&\beta I\end{bmatrix}$. Therefore, the Jacobian of $T$ at $(x^{*},\epsilon^{*})$ is given by
$DT(x^{*},\epsilon^{*})=$ $\ \begin{bmatrix}(1-\alpha)I+\alpha(I-\tfrac{1}{\beta}D_{x}g(x^{*},\epsilon^{*}))\ &D_{\epsilon}g(x^{*},\epsilon^{*})\\ \bm{0}\ &(1-\alpha(1-\mu))I\end{bmatrix}$
$=$ $\ \begin{bmatrix}I-\frac{\alpha}{\beta}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})&I-\frac{\alpha}{\beta}\nabla^{2}_{\mathcal{I}^{*}\mathcal{J}^{*}}f(x^{*})&D_{\epsilon}g(x^{*},\epsilon^{*})\\ \bm{0}&(1-\alpha)I&\bm{0}\\ \bm{0}&\bm{0}&(1-\alpha(1-\mu))I\end{bmatrix}$
@@
If $\alpha$ is selected to satisfy $01-\frac{\alpha}{\beta}\lambda_{\max}(\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))>$ 0 and by the strict saddle property $\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})$ does not have 0 eigenvalue, we know $DT(x^{*},\epsilon^{*})$ is invertible. ∎
@@
(Lipeomorphism) Suppose Assumption 1-3 hold. Let $L_{r}:=r^{\prime\prime}(|(r^{\prime})^{-1}(C))$ if $r^{\prime}(|0^{+}|)=\infty$, where $C$ defined in Proposition 10, and $L_{r}:=r^{\prime\prime}(|0^{+}|)$ if $r^{\prime}(|0^{+}|)<\infty$. Then $T$ is invertible, and $T$ and $T^{-1}$ is Lipschitz continuous (i.e., $T$ is a lipeomorphism) when $10<\alpha<$ 1, $\beta>\alpha L_{f}/2$ and $\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))<1$.
@@
If $r^{\prime}(|0+|)<\infty$, $w(x_{i},\epsilon_{i})=r^{\prime}(|x_{i}|,\epsilon_{i})$ is Lipschitz continuous in $\mathcal{L}(F^{0})$.
@@
$u_{i}(x_{i},\epsilon_{i})=\begin{cases}w(x_{i},\epsilon_{i})=r^{\prime}(|x_{i}|+\epsilon_{i})&\text{ if }r^{\prime}(|0+|)<\infty\\ \min(w(\underline{x},0),w(x_{i},\epsilon_{i}))=\min(r^{\prime}(\underline{x}),r^{\prime}(|x_{i}|+\epsilon_{i}))&\text{ if }r^{\prime}(|0+|)=\infty.\end{cases}$
@@
${\|}S^{x}(\hat{x},\hat{\epsilon})-S^{x}(\tilde{x},\tilde{\epsilon}){\|}=$ $\ {\|}\mathbb{S}_{\lambda\hat{u}/\beta}(\hat{x}-\tfrac{1}{\beta}\nabla f(\hat{x}))-\mathbb{S}_{\lambda\tilde{u}/\beta}(\tilde{x}-\tfrac{1}{\beta}\nabla f(\tilde{x})){\|}$
@@
where the first inequality is by Lemma A.3 and the second is by (19) and (30). Therefore, $\|\alpha(\hat{x}-S^{x}(\hat{x},\hat{\epsilon}))-\alpha(\tilde{x}-S^{x}(\tilde{x},\tilde{\epsilon}))\|\leq\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))$ holds for any $\hat{y},\tilde{y}\in\mathcal{L}(F^{0})$ and $\hat{\epsilon},\tilde{\epsilon}\in\mathbb{R}^{n}_{+}$.
Notice that $T=\begin{bmatrix}x\\ \epsilon\end{bmatrix}+\alpha(\begin{bmatrix}S^{x}(x,\epsilon)\\ \mu\epsilon\end{bmatrix}-\begin{bmatrix}x\\ \epsilon\end{bmatrix})$, Since $\begin{bmatrix}S^{x}(x,\epsilon)\\ \mu\epsilon\end{bmatrix}$ is Lipschitz with constant $\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))<1$. Applying [23, Lemma 2.13], we directly derive that $T=I+\alpha(S-I)$ is invertible and a lipeomorphism. ∎
@@
(Global Escape) Consider problem (P) with strict saddle property and the ${\rm DIRL}_{1}$ algorithm with Assumption 1-3. Let the condition of Proposition 14 holds and $\alpha<\beta/\rho$. Let $\{(x^{k},\epsilon^{k})\}$ be the sequence generated by the Algorithm 10. Then, the set of initial points attracted by strict saddle points has zero Lebesgue measure.